# Statistical testing of random number generators and their improvement using randomness extraction

### v1

@@ V.1 Overview
    We collect the output of each RNG that we tested in the previous section. We call this the initial output.
@@
    We apply different post-processing methods, or randomness extractors, to this initial output to produce a new, processed, output. Each time, we precisely define and explain the underlying assumptions of the used extractors required for the extraction method to be successful. These different sets of assumptions, for each extraction method, can be compared with each other and form the different post-processing levels.
@@
    We analyse the new, processed output with our $\mathsf{STE}$ to determine whether each extraction method had an impact from a statistical perspective. We also compare the results obtained using the different post-processing methods, for each RNG.
@@
V.2 Randomness extraction methods
@@
We now describe the different post-processing levels we consider in this work, i.e. the types of randomness extractors that we will use to improve the different RNGs. We consider four classes of randomness extractor, which form the different levels, each with increasingly elaborate implementations:
  • Level 1
@@
    Deterministic extractors - this class of extractors requires certain properties of the initial output’s distribution to hold, beyond just a min-entropy assumption. An example is the seminal Von Neumann extractor von1963various , which works if every bit of the initial output is identically and independently generated (although a sufficient condition is that the input forms an exchangeable sequence). In practice, assumptions of this type are often difficult to justify and hard to control.
  • Level 2
@@
    Seeded extractors - these extractors require a second string, called a seed, of independent and (near-)perfectly random bits as the resource. This seed needs to be carefully generated, and can lead to problems if, for example, it is not generated independently of the initial output of the RNG [^8] This could happen, for example, if the seed is generated whilst sharing the same environment as the RNG or by an adversary. or if it has poor statistical properties. At a fundamental level, seeded extractors are unsatisfying as there is a circularity in having to generate near-perfect randomness as a resource to build an RNG.
  • Level 3
@@
    Two-source extractors - these extractors are a generalisation of seeded extractors in which the assumptions on the seed are relaxed. Namely, the second, additional source of randomness (previously the seed) now only needs to have some known min-entropy and be independent from the initial output. Moreover, the independence condition can also be relaxed, for example allowing coordination, cross influence or bounded mutual information with respect to the input ball2022randomness or independence only in the sense of a Markov chain arnon2015quantum .
  • Level 4
@@
    Physical device-independent extractors - the last class that we consider are extractors requiring special additional hardware, providing the second randomness source needed in level 3 whilst making only minimal assumptions [^9] For example, that information cannot travel faster than the speed of light.. This is made possible by a particular type of interactive proof system in which quantum hardware can be verified to perform as promised, as opposed to having to rely on modelling the physical process as would be done normally. This ‘black box’ verification gives a guaranteed lower bound on the min-entropy of the output, which can then be used together with the RNG’s initial output in a two-source extractor as in level 3. These physical extractors are referred to, in the quantum information science community, as device-independent randomness amplification protocols and have no classical analogue. With today’s technology such extractors require making a few additional implementation assumptions (to the minimal ones). We come back in detail to physical extractors in Section V.8.
@@
When a second bit string of randomness is required (levels 2 and 3), we use the NIST Randomness Beacon kelsey2019reference . For physical randomness extraction (level 4), we use a semi-device-independent randomness amplification protocol that is an adaptation of foreman2020practical , which we describe in Section V.8. All the algorithms for extraction used in this work are from the software library Cryptomite cryptomite , which can be found at https://github.com/CQCL/cryptomite.
@@
Figure 3: Illustration of the set of sources, or input distributions, that can be successfully extracted from by different randomness extraction methods. (Right) weak input distributions and (Left) second input, or weak seed, distributions. Deterministic extractors (level 1) require additional properties on the weak input, but do not need a second input source. Seeded extractor (level 2) relax the need for additional properties of the weak input and extract from sources with min-entropy only, at the cost of requiring a second string of (near-)perfect randomness. Two-source extractors (level 3) relax the assumptions of seeded ones to a second source that also has min-entropy only. Physical extractors (level 4, not on the figure) requires special quantum hardware, which effectively provides the second input with a device-independent lower bound on the min-entropy, requiring minimal added assumptions.
V.3 Results overview
@@
We now present the main results of statistical testing the different post-processing methods in Figure 4, with more details and tables in the following sections. As stated before, we expect that an RNG producing near-perfect random numbers fails less than 7.5 of the 4600 tests it is subject to, on average, when testing $5\times 10$Gbit files [^10] This number is the expected amount of type 2 error, i.e. the expected maximum number of failed tests, given that the underlying distribution is indistinguishable from uniform. Note that we implicitly assume that each statistical test is independent.. This is the criterion we use to call randomness generation successful from a statistical perspective (green highlighted area).
Refer to captionRefer to caption
@@
Figure 4: The above plots show (left) the number of statistical tests failed and (right) failed and suspicious for each initial RNG at each post-processing level. The $x$ axis indicates the level, with step 0 being the initial RNG with no additional post-processing and steps 1-4 are deterministic, seeded, two-source and physical extraction, respectively. The $y$ axis is the number of statistical tests failed (left) or failed and suspicious (right), out of 4600, using a logarithmic scale: for $f$ failed or failed and suspicious tests, $y=\log_{2}(f+1)$. The shaded region on the left plot illustrates the successful region, whereby the RNG fails less than 7.5 tests, and the white region illustrates the ‘unacceptable’ region, in which, with high probability, near-perfect randomness is not produced. We note that we are unable to use the 32-bit LFSR at level 4 because of its low initial estimated min-entropy rate, $\alpha_{\mathsf{RNG}}$, as detailed and evaluated in Section IV.
@@
Ideally, the results would reflect the different levels of the post-processing and the validity of the assumptions these imply. Our results in Figure 4 tell a mixed story.
@@
  • For the RNGs that fail the tests when unprocessed, we observe that additional post-processing indeed improves the quality of the initial output. Considering the LFSR, for example, any extraction method higher than level 1 applied to the initial output produced a processed output which passed the numerical tests well. IDQ’s device, as a second example, is significantly improved already with level 1 of extraction, but only gives successful results when higher levels are applied.
@@
  • Although level 3 is strictly a relaxation of the assumptions made at level 2, we were unable to observe a difference in the numerical results. This is because level 2, from a statistical perspective, seems to be giving results that are already successful. Moreover, we are unable to distinguish between level 2, 3 and 4. We interpret this as another illustration of the difference between statistical and cryptographic randomness, in which weaker assumptions are desirable even if no statistical advantage can be witnessed from the user’s perspective. It is also likely that, in order to give nontrivial examples of step 2 failing, one would need to generate the seed in a manner that is either significantly biased or correlated to the RNG (both of which could happen in practice).
@@
  • All our implementations above level 1 give successful numerical test results on the three RNGs that we tested. In particular, from a statistical perspective, this means that a poor PRNG (here the 32-bit LFSR) can be concatenated with an extractor to form a good PRNG.
V.4 Implementations of the post-processing methods
@@
We now describe how we implemented the post-processing, i.e. different extractors in our levels, together with the parameter choices and compromises we made. For the post-processing algorithms, we use the randomness extractors publicly available from the software library Cryptomite cryptomite . In order to test the randomness quality at each step we generate $5\times 10$Gbit test files of processed output and perform statistical testing with the ‘all’ setting (the most intense) in the $\mathsf{STE}$.
All randomness post-processing and statistical tests were run on a Dell Precision 7540 personal laptop with 16GB of RAM and a 2.3GHz Intel i9 processor, using the Ubuntu 20.04 operating system. We state all input and output sizes and give detailed descriptions of each test setting and implementation of each level with the parameter choices, so that all results can be reproduced.
For each level, we choose the parameters of the different extractors such that, in theory, the processed output is $\epsilon_{\mathsf{total}}$-perfectly random (see Definition 5), with $\epsilon_{\mathsf{total}}\leq 2^{-32}\approx 10^{-10}$.
V.5 Level 1: Deterministic extraction
@@
Definition 7 (Deterministic randomness extractor).
A deterministic randomness extractor is a function
@@
such that, for random variables $X\in\{0,1\}^{n}$ with specific properties shaltiel2011introduction ,
$\Delta(\mathsf{Ext_{d}}(X),\mathrm{U}_{m})\leq\epsilon$ (8)
@@
In words, a deterministic extractor is a function that maps random variables $X$ with specific characteristics, to a new variable $\mathsf{Ext_{d}}(X)$ that is near-perfectly random. Note that the properties of $X$ required depend on the specific extractor – for example that all the bits in $X$ are I.I.D.. In practice, those properties are hard (or even impossible) to verify and it is preferable to make a claim about the min-entropy only.
The implementation of the deterministic extraction set-up is shown in Figure 5.
Figure 5: The level 1 of our post-processing methods is performed by using a deterministic extractor, namely the Von Neumann extractor, on the initial output of the RNG.
@@
We use the Von Neumann extractor von1963various to extract from the initial output $X\in\{0,1\}^{n}$ of the RNG, with the implementation from cryptomite . This extractor requires that all two subsequent input bits have a fixed bias, i.e. for bits $X_{2i},X_{2i+1}\in\{0,1\}$ with $i=1,\ldots,\lfloor\frac{n}{2}\rfloor$ and $p_{i}\in(0,1)$, we require that
$\Pr(X_{2i}=0)=\Pr(X_{2i+1}=0)=p_{i}$ (9)
@@
The Von Neumann extractor works by grouping subsequent bits in pairs, and outputting the first (or second) bit only when the bits in the pair are different, giving an output length of $m\approx p(1-p)$ (if the bias is fixed $p_{i}=p$ for all $i$) and $\epsilon=0$, i.e. perfect randomness at the output.
V.5.1 Results
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
32-bit LFSR 25 10 (5) 5 18 76 106 724 (413)
RDSEED 0 0 (2) 0 0 0 1 0 (2)
IDQ Quantis 4 0 (1) 0 0 0 3 0 (3)
Table 5: This table gives the sum of tests failed for $5\times 10$Gbit samples from each RNG, after deterministic extraction using the Von Neumann extractor. Due to the 32-bit LFSR failing PractRand quickly, only 635 tests were conducted (instead of the full 4600) so we re-scale these results. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses. Full results can be found in Appendix C.
The statistical test results show that:
@@ V.5.1 Results
  • There is an observed improvement, for both the LFSR and IDQ Quantis, compared with the results of the initial RNG testing (Table 3), although they are still not successful.
@@
  • The number of NIST statistical tests failed by both the LFSR and IDQ Quantis increases when moving from no post-processing to deterministic post-processing. This could happen for many reasons, including that the RNGs have a specific bias structure that is incompatible (and indeed amplified) by the Von Neumann extractor or that there are some fundamental issues with the NIST tests (as suggested in burciu2019systematic ; hamano2007correction ; kowalska2022revision ; marton2015interpretation ).
V.6 Level 2: Seeded extraction
The properties required for deterministic extraction (level 1) from an RNG to be successful are hard to justify in practice. Seeded extraction requires only that the initial RNG output has a min-entropy guarantee, i.e. that it is only somewhat random – a much weaker requirement on the initial RNGs output, making it more realistic in practice. The cost for this weaker requirement is that a second, independent and (near-)perfectly random string (a seed) now needs to be provided.
Definition 8 (Seeded randomness extractor).
@@ V.6 Level 2: Seeded extraction
A seeded randomness extractor is a function $\mathsf{Ext_{s}}:\{0,1\}^{n}\times\{0,1\}^{d}\rightarrow\{0,1\}^{m}$ such that, for a random variable $X\in\{0,1\}^{n}$ with min-entropy $\mathrm{H}_{\infty}(X)\geq k$, and seed $S\in\{0,1\}^{d}$ with min-entropy $\mathrm{H}_{\infty}(S)=d$ (i.e. $S$ is perfectly random) then,
$\Delta(\mathsf{Ext_{s}}(X,S),\mathrm{U}_{m})\leq\epsilon$ (10)
@@
A seeded extractor can be understood as a randomized function that maps a weakly random variable $X$ to a new variable $\mathsf{Ext_{s}}(X,S)$ that is (near-)perfectly random. Note that the seed may be $\epsilon_{s}$-perfect only, with additive error in the statistical distance above, i.e. $\epsilon\rightarrow\epsilon+\epsilon_{s}$ (see, for example, Appendix A from frauchiger4547true for a proof). Seeded extractors are a special case of a two-source extractor, which we define later in Definition 10.
Definition 9 (Strong seeded extractor).
A strong seeded randomness extractor is a function $\mathsf{Ext_{s}}:\{0,1\}^{n}\times\{0,1\}^{d}\rightarrow\{0,1\}^{m}$ such that
$\Delta([\mathsf{Ext_{s}}(X,S),S],[\mathrm{U}_{m},S])\leq\epsilon$ (11)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of random variables and $\mathrm{U}_{m}$ is the uniform variable on $\{0,1\}^{m}$.
A strong seeded extractor is a randomized function that gives a (near-)uniform output, even when conditioned on the seed $S$ (the output is therefore independent of the seed). This has some interesting consequences, which we exploit to generate the large amounts of processed output needed for statistical testing. Specifically, $S$ can be re-used with different weak input random variables, allowing a single seed to be used in many extraction rounds. The step of seeded extraction (implemented using a strong seeded extractor) is shown in Figure 6. The initial output from the RNG is split into blocks $X_{i}$ for $i=1,\ldots,n$ with a promise on each block’s min-entropy (Definition 2).
@@
This step can be implemented with the Circulant cryptomite , Dodis et al. dodis2004improved , Toeplitz krawczyk1994lfsr , and Trevisan trevisan1999construction extractors from Cryptomite, as they can all be used as strong seeded extractors. Among these extractors, Circulant offers the best trade-off between security parameters and computational complexity, and is therefore the one we chose. The Circulant extractor requires that the seed length is the input length plus one, and that the seed length is a prime. We set the seed length $|S|$ and RNG input block lengths $|X_{i}|$ to $|S|=|X_{i}|+1=10007$. Note that, using Circulant allows to generate cryptographic randomness even against an adversary able to store (and process) side-information in quantum systems without changing the extraction algorithm – i.e. the extractor is quantum-proof, see cryptomite for details.
To generate the seed $S$, we use the NIST Randomness Beacon, which is a public source of randomness produced by the US Government agency (NIST) mixing different randomness sources together, including chaotic classical and quantum processes kelsey2019reference . The min-entropy $k_{i}^{\mathsf{RNG}}$ for each block $X_{i}$ is $k_{i}^{\mathsf{RNG}}=\alpha_{\mathsf{RNG}}|X_{i}|$, where $\alpha_{\mathsf{RNG}}$ is a lower bound on the min-entropy per bit for each initial RNG block of outputs $X_{i}$, with probability $\epsilon_{\mathsf{est}}<2^{-32}$ (as found in Equation 19 of Section IV). The output length after extraction, $m$, is then roughly $m\approx k_{i}^{\mathsf{RNG}}$.
In order to generate the required $5\times 10$Gbit’s of processed output, the Circulant extractor is used multiple times on different initial output blocks $X_{i}$ with the same seed. The extractor’s outputs are then concatenated together until a final output, $\mathsf{Output}$, of sufficient size is generated. The $\mathsf{Output}$ is given by:
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of random variables. Each extraction round, which we index $i$, has an associated error $\epsilon_{\mathsf{ext_{i}}}$ and we choose the total security parameter to be $\epsilon_{\mathsf{total}}\leq 2^{-32}$ – namely, everything is chosen so that $\epsilon_{\mathsf{total}}=\epsilon_{\mathsf{est}}+\sum_{j=1}^{n}\epsilon_{\mathsf{ext_{j}}}\leq 2^{-32}$. This derivation for $\epsilon_{\mathsf{total}}$, specifically that the composed output error is the sum of each of the individual extractor errors, can be found in cryptomite .
V.6.1 Results
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
32-bit LFSR 0 0 (3) 0 0 0 0 0 (6)
RDSEED 0 0 (7) 0 0 0 0 0 (7)
IDQ Quantis 0 0 (2) 0 0 0 2 0 (5)
Table 6: This table gives the sum of tests failed for $5\times 10$Gbit samples from each RNG, after a strong seeded extractor has been applied to its initial output. The seed is generated using the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right. Full results can be found in Appendix D.
The observations we draw from the results in Table 6 are the following:
  • The statistical test results show a significant improvement on the results using deterministic extraction, see Section V.5. In particular, now all RNGs have been successfully post-processed from a statistical perspective.

### v2

@@ 5 A Variety of Post-Processing Methods
    We collect the output of each RNG that was tested in the previous section. We call this the initial output.
@@
    We apply different post-processing methods, or randomness extractors, to this initial output to produce a new, processed output. Each time, we precisely define and explain the underlying assumptions of the used extractors required for the extraction method to be successful. These different sets of assumptions, for each extraction method, can be compared with each other and form the different post-processing levels.
@@
    We analyse the new, processed output with our $\mathsf{STE}$ to determine whether each extraction method had an impact from a statistical perspective. We also compare the results obtained using the different post-processing methods for each RNG.
@@
5.1 Randomness Extraction Methods
@@
We now describe the different post-processing levels that we consider in this work, i.e., the types of randomness extractors that we will use to improve the different RNGs. We consider four classes of randomness extractor, which form the different levels, each with increasingly elaborate implementations.
@@
  • Level 1: Deterministic extractors—This class of extractors requires certain properties of the initial output’s distribution to hold, beyond just a min-entropy assumption. An example is the seminal Von Neumann extractor Von Neumann (1963), which works if every bit of the initial output is identically and independently generated (although a sufficient condition is that the input forms an exchangeable sequence). In practice, assumptions of this type are difficult to justify and to control.
@@
  • Level 2: Seeded extractors—These extractors require a second string, called a seed, of independent and (near-)perfectly random bits as the resource. This seed needs to be carefully generated and can lead to problems if, for example, it is not generated independently of the initial output of the RNG (this could happen, for example, if the seed is generated whilst sharing the same environment as the RNG or by an adversary) or if it has poor statistical properties. At a fundamental level, seeded extractors are unsatisfying as there is circularity in having to generate near-perfect randomness as a resource to build an RNG.
@@
  • Level 3: Two-source extractors—These extractors are a generalisation of seeded extractors in which the assumptions on the seed are relaxed. Namely, the second, additional source of randomness (previously the seed) now only needs to have some known min-entropy and be independent of the initial output. Moreover, the independence condition can also be relaxed—for example, allowing coordination, cross-influence, or bounded mutual information with respect to the input Ball et al. (2022) or independence only in the sense of a Markov chain Arnon-Friedman et al. (2015).
@@
  • Level 4: Physical device-independent extractors—The last class that we consider are extractors requiring special additional hardware, providing the second randomness source needed in level 3 whilst making only minimal assumptions (for example, that information cannot travel faster than the speed of light). This is made possible by a particular type of interactive proof system in which quantum hardware can be verified to perform as promised, as opposed to having to rely on modelling the physical process, as would be done normally. This ‘black box’ verification gives a guaranteed lower bound on the min-entropy of the output, which can then be used together with the RNG’s initial output in a two-source extractor as in level 3. These physical extractors are referred to, in the quantum information science community, as device-independent randomness amplification protocols and have no classical analogue. With today’s technology, such extractors require making additional implementation assumptions (to the minimal ones). We return in detail to physical extractors in Section 5.2.4.
@@
When a second bit string of randomness is required (levels 2 and 3), we use the NIST Randomness Beacon Kelsey et al. (2019). For physical randomness extraction (level 4), we use a semi-device-independent randomness amplification protocol that is an adaptation of Foreman et al. (2023), which we describe in Section 5.2.4. All the algorithms for extraction used in this work are from the software library Cryptomite Foreman et al. (2024), which can be found at https://github.com/CQCL/cryptomite (accessed on 28 November 2024).
@@
Figure 3: Illustration of the set of sources, or input distributions, that can be successfully extracted from by different randomness extraction methods. (Right) weak input distributions and (Left) second input, or weak seed, distributions. Deterministic extractors (level 1) require additional properties on the weak input but do not need a second input source. Seeded extractors (level 2) relax the need for additional properties of the weak input and extract from sources with min-entropy only, at the cost
of requiring a second string of (near-)perfect randomness. Two-source extractors (level 3) relax the assumptions of seeded ones to a second source that also has min-entropy only. Physical extractors (level 4, not in the figure) require special quantum hardware, which effectively provides the second input with a device-independent lower bound on the min-entropy, requiring minimal added assumptions.
5.1.1 Results Overview
@@
We now present the main results of the statistical testing of the different post-processing methods in Figure 4, with more details and tables in the following sections. As stated before, we expect that an RNG producing near-perfect random numbers fails less than 7.5 of the 4600 tests that it is subject to, on average, when testing $5\times 10$ Gbit files. (This number is the expected amount of type 2 error, i.e., the expected maximum number of failed tests, given that the underlying distribution is indistinguishable from uniform. Note that we implicitly assume that each statistical test is independent). This is the criterion that we use to call randomness generation successful from a statistical perspective (green highlighted area).
@@
Figure 4: The above plots show (left) the number of statistical tests failed and (right) failed and suspicious for each initial RNG at each post-processing level. The $x$ axis indicates the level, with step 0 being the initial RNG with no additional post-processing, and steps 1–4 are deterministic, seeded, two-source, and physical extraction, respectively. The $y$ axis is the number of statistical tests failed (left) or failed and suspicious (right), out of 4600, using a logarithmic scale: for $f$ failed or failed and suspicious tests, $y=\log_{2}(f+1)$. The shaded region in the left plot illustrates the successful region, whereby the RNG fails less than 7.5 tests, and the white region illustrates the ‘unacceptable’ region, in which, with high probability, near-perfect randomness is not produced. We note that we are unable to use the 32-bit LFSR at level 4 because of its low initial estimated min-entropy rate, $\alpha_{\mathsf{RNG}}$, as detailed and evaluated in Section 4.
@@
Ideally, the results would reflect the different levels of post-processing and the validity of the assumptions that these imply. Our results in Figure 4 tell a mixed story.
@@
  • For the RNGs that fail the tests when unprocessed, we observe that additional post-processing indeed improves the quality of the initial output. Considering the LFSR, for example, any extraction method higher than level 1 applied to the initial output produced a processed output that passed the numerical tests well. IDQ’s device, as a second example, is significantly improved already with level 1 of extraction, but only gives successful results when higher levels are applied.
@@
  • Although level 3 is strictly a relaxation of the assumptions made at level 2, we were unable to observe a difference in the numerical results. This is because level 2, from a statistical perspective, seems to be giving results that are already successful. Moreover, we are unable to distinguish between levels 2, 3, and 4. We interpret this as another illustration of the difference between statistical and cryptographic randomness, in which weaker assumptions are desirable even if no statistical advantage can be witnessed from the user’s perspective. It is also likely that, in order to give non-trivial examples of step 2 failing, one would need to generate the seed in a manner that is either significantly biased or correlated to the RNG (both of which could happen in practice).
@@
  • All our implementations above level 1 gave successful numerical test results on the three RNGs that we tested. In particular, from a statistical perspective, this means that a poor PRNG (here, the 32-bit LFSR) can be concatenated with an extractor to form a good PRNG.
5.2 Implementations of the Post-Processing Methods
@@
We now describe how we implemented the post-processing, i.e., different extractors in our levels, together with the parameter choices and compromises that we made. For the post-processing algorithms, we used the randomness extractors publicly available from the software library Cryptomite Foreman et al. (2024). To assess the randomness quality at each step, we generated $5\times 10$ Gbit test files of the processed output and performed statistical testing using the ‘all’ setting (the most intense) in the $\mathsf{STE}$. All randomness post-processing and statistical tests were run on a Dell Precision 7540 personal laptop with 16 GB of RAM and a 2.3 GHz Intel i9 processor, using the Ubuntu 20.04 operating system. We state all input and output sizes and give detailed descriptions of each test setting and implementation of each level with the parameter choices, so that all results can be reproduced. For each level, we chose the parameters of the different extractors such that, in theory, the processed output was $\epsilon_{\mathsf{total}}$-perfectly random (see Figure 2), with $\epsilon_{\mathsf{total}}\leq 2^{-32}\approx 10^{-10}$.
5.2.1 Level 1: Deterministic Extraction
@@
{Definition}
[Deterministic randomness extractor] A deterministic randomness extractor is a function
@@
such that, for random variables $X\in\{0,1\}^{n}$ with specific properties Shaltiel (2011),
$\Delta(\mathsf{Ext_{d}}(X),\mathrm{U}_{m})\leq\epsilon\ ,$ (7)
@@
In other words, a deterministic extractor is a function that maps random variables $X$ with specific characteristics to a new variable $\mathsf{Ext_{d}}(X)$ that is near-perfectly random. Note that the properties of $X$ required depend on the specific extractor—for example, that all bits in $X$ are I.I.D.
@@
The implementation of the deterministic extraction set-up is shown in Figure 5. We use the Von Neumann extractor Von Neumann (1963) to extract from the initial output $X\in\{0,1\}^{n}$ of the RNG, with the implementation from Foreman et al. (2024). This extractor requires that all two subsequent input bits have a fixed bias, i.e., for bits $X_{2i},X_{2i+1}\in\{0,1\}$ with $i=1,\ldots,\lfloor\frac{n}{2}\rfloor$ and $p_{i}\in(0,1)$, we require that
$\Pr(X_{2i}=0)=\Pr(X_{2i+1}=0)=p_{i}\ .$ (8)
@@
The Von Neumann extractor works by grouping subsequent bits in pairs and outputting the first (or second) bit only when the bits in the pair are different, giving an output length of $m\approx p(1-p)$ (if the bias is fixed, $p_{i}=p$ for all $i$) and $\epsilon=0$, i.e., perfect randomness at the output.
Figure 5: Here, level 1 of our post-processing methods is performed by using a deterministic extractor, namely the Von Neumann extractor, on the initial output of the RNG.
The statistical test results in Table 5 show the following.
@@
  • Both the LFSR and IDQ Quantis show improvements compared to the initial RNG testing results (Table 3), although they still do not pass overall.
@@
  • The number of NIST statistical test failures for both the LFSR and IDQ Quantis increases when moving from no post-processing to deterministic post-processing. This may be due to specific biases in the RNGs that are incompatible with, or even amplified by, the Von Neumann extractor (e.g., successive bits are not independent) or fundamental issues with the NIST tests, as suggested in Burciu and Simion (2019); Hamano and Kaneko (2007); Kowalska et al. (2022); Marton and Suciu (2015).
Table 5: This table gives the sum of statistical tests failed for $5\times 10$ Gbit samples from each RNG, after deterministic extraction using the Von Neumann extractor. Due to the 32-bit LFSR failing PractRand quickly, only 635 tests were conducted (instead of the full 4600), so we rescale these results. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses. Full results can be found in Appendix C.
{adjustwidth}
[…]
Interestingly, applying the Von Neumann extractor to an LFSR results in a stream cipher known as the self-shrinking generator, which has been studied for cryptographic use Meier and Staffelbach (1994). Although the self-shrinking generator fails fewer tests compared to the unprocessed LFSR, a substantial number of failures remain.
As noted at the start of this section, a deterministic extractor can produce a near-perfectly random output if the input source satisfies certain specific properties. However, these properties are often difficult or even impossible to verify in practice. As a result, it is more practical to base claims solely on the min-entropy of the source. In Santha and Vazirani (1986), it was demonstrated that deterministic extraction from a source characterised solely by min-entropy is impossible. Such sources require additional, independent randomness to enable extraction.
5.2.2 Level 2: Seeded Extraction
Seeded extraction requires only a min-entropy guarantee for the initial RNG output but comes at the cost of needing a second, independent and (near-)perfectly random input (the seed) to enable extraction.
{Definition}
@@ 5.2.2 Level 2: Seeded Extraction
[Seeded randomness extractor] A seeded randomness extractor is a function $\mathsf{Ext_{s}}:\{0,1\}^{n}\times\{0,1\}^{d}\rightarrow\{0,1\}^{m}$ such that, for a random variable $X\in\{0,1\}^{n}$ with min-entropy $\mathrm{H}_{\infty}(X)\geq k$, and seed $S\in\{0,1\}^{d}$ with min-entropy $\mathrm{H}_{\infty}(S)=d$ (i.e., $S$ is perfectly random),
$\Delta(\mathsf{Ext_{s}}(X,S),\mathrm{U}_{m})\leq\epsilon\ ,$ (9)
@@
A seeded extractor can be understood as a randomised function that maps a weakly random variable $X$ to a new variable $\mathsf{Ext_{s}}(X,S)$ that is (near-)perfectly random. Note that the seed may be $\epsilon_{s}$-perfect only, with an additive error in the statistical distance above, i.e., $\epsilon\rightarrow\epsilon+\epsilon_{s}$ (see, for example, Appendix A from Frauchiger et al. for proof). Seeded extractors are a special case of two-source extractors, which we define later in Section 5.2.3.
{Definition}
[Strong seeded extractor] A strong seeded randomness extractor is a function $\mathsf{Ext_{s}}:\{0,1\}^{n}\times\{0,1\}^{d}\rightarrow\{0,1\}^{m}$ such that
$\Delta([\mathsf{Ext_{s}}(X,S),S],[\mathrm{U}_{m},S])\leq\epsilon\ ,$ (10)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of random variables and $\mathrm{U}_{m}$ is the uniform variable on $\{0,1\}^{m}$. A strong seeded extractor is a randomised function that gives a (near-)uniform output, even when conditioned on the seed $S$ (the output is therefore independent of the seed). This has some interesting consequences, which we exploit to generate the large amounts of processed output needed for statistical testing. Specifically, $S$ can be reused with different weak input random variables, allowing a single seed to be used in many extraction rounds. The set-up for seeded extraction (implemented using a strong seeded extractor) is shown in Figure 6. The initial output from the RNG is split into blocks $X_{i}$ for $i=1,\ldots,n$ with a promise on each block’s min-entropy (Figure 2).
@@
This step can be implemented with the Circulant Foreman et al. (2024), Dodis et al. Dodis et al. (2004), Toeplitz Krawczyk (1994), and Trevisan Trevisan (1999) extractors from Cryptomite, as they can all be used as strong seeded extractors. Among these extractors, Circulant offers the best trade-off between security parameters and computational complexity and is therefore the one that we chose. The Circulant extractor requires that the seed length is the input length plus one and that the seed length is a prime. We set the seed length $|S|$ and RNG input block lengths $|X_{i}|$ to $|S|=|X_{i}|+1$ = 10,007. Note that using Circulant allows us to generate cryptographic randomness even against an adversary able to store (and process) side information in quantum systems without changing the extraction algorithm, i.e., the extractor is quantum-proof; see Foreman et al. (2024) for details.
To generate the seed $S$, we use the NIST Randomness Beacon, which is a public source of randomness produced by the US Government agency (NIST), mixing different randomness sources together, including chaotic classical and quantum processes Kelsey et al. (2019). The min-entropy $k_{i}^{\mathsf{RNG}}$ for each block $X_{i}$ is $k_{i}^{\mathsf{RNG}}=\alpha_{\mathsf{RNG}}|X_{i}|$, where $\alpha_{\mathsf{RNG}}$ is a lower bound on the min-entropy per bit for each initial RNG block of outputs $X_{i}$, with probability $\epsilon_{\mathsf{est}}<2^{-32}$ (as found in Equation 18 of Section B.2.1). The output length after extraction, $m$, is then roughly $m\approx k_{i}^{\mathsf{RNG}}$.
In order to generate the required $5\times 10$ Gbits of processed output, the Circulant extractor is used multiple times on different initial output blocks $X_{i}$ with the same seed. The extractor’s outputs are then concatenated together until a final output, $\mathsf{Output}$, of sufficient size is generated. The $\mathsf{Output}$ is given by
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of random variables. Each extraction round, which we index $i$, has an associated error $\epsilon_{\mathsf{ext_{i}}}$, and we choose the total security parameter to be $\epsilon_{\mathsf{total}}\leq 2^{-32}$—namely, everything is chosen so that $\epsilon_{\mathsf{total}}=\epsilon_{\mathsf{est}}+\sum_{j=1}^{n}\epsilon_{\mathsf{ext_{j}}}\leq 2^{-32}$. This derivation for $\epsilon_{\mathsf{total}}$, specifically that the composed output error is the sum of each of the individual extractor errors, can be found in Foreman et al. (2024).
The observations that we draw from the results in Table 6 are the following.
  • The statistical test results show a significant improvement in the results using deterministic extraction; see Section 5.2.1. In particular, all RNGs have been successfully post-processed from a statistical perspective.